# TomoGraphView: 3D Medical Image Classification with Omnidirectional Slice Representations and Graph Neural Networks

### v1

@@ 3 Methodology
In this work, we present TomoGraphView, a novel framework that preserves the spatial structure of volumetric medical imaging data when decomposed into 2D slices compatible with 2D feature encoders. The name TomoGraphView encapsulates its core components: Tomo refers to tomography data, Graph highlights the GNN-based feature aggregation, and View emphasizes the integration of multi-view information from omnidirectional perspectives. In general, our framework is built upon two key components: omnidirectional volume slicing and spherical mesh–based graph topology construction with corresponding GNN-based feature aggregation. A visual summary of the method is depicted in Figure 1. The following sections detail each component of TomoGraphView. We first describe omnidirectional volume slicing in Section 3.1 and relate it to conventional slicing approaches. Section 3.2 outlines the encoding of the obtained omnidirectional 22D cross-sectional slices into 11D feature representations. In Section 3.3, we introduce the spherical mesh–based graph topology construction, followed by Section 3.4, which describes the graph representation learning process used for downstream classification.
@@ 3.1 Omnidirectional Volume Slicing
Figure 2: Visual representation of different volume slicing strategies and resulting slices for a synthetically generated input volume depicting an elongated shape diagonally within the volume. Volume slicing strategies include 22D-axial and 22D-axial+ (left), 2.52.5D and 2.52.5D+ (middle), and our proposed omnidirectional volume slicing (right). Green slices represent slices extracted from the axial direction, red from the sagittal direction, and blue from the coronal direction, whereas yellow shows out-of-plane slices as employed within our omnidirectional volume slicing strategy.
@@
In addition to utilizing information from canonical planes, we propose to extract slices from non-canonical orientations to obtain omnidirectional slice representations, thereby enabling a richer characterization of volumetric structures that are not naturally aligned with the canonical axes. To obtain $N$ omnidirectional slice representations, we visually embed the volume within a bounding sphere $\mathcal{S}$, centered at the origin, which is formally described as
@@
where $r$ is the radius such that the sphere fully contains the 33D volumetric scan as illustrated in Figure 3 (step I). Next, as we intend to make use of both, canonical and non-canonical perspectives, we first of all fix three points on the surface of the sphere representing the canonical view planes (i.e., axial, coronal, and sagittal) with their known coordinates $(r,0,0),(0,r,0)$ and $(0,0,r)$, as illustrated in Figure 3 (step I).
We then optimize the location of the remaining (N3)(N-3) points, representing non-canonical view planes, such that all points on the surface are approximately equally distributed, as illustrated in Figure 3 (step II). To this end, we make use of the repulsion-based optimization outlined in Algorithm 1, which is inspired by the Thomson problem [thomson1904xxiv]. Therein, points are interpreted as identical charged particles constrained to lie on the spherical surface, and their positions are optimized by minimizing the system’s total electrostatic energy. Formally, given $N$ points $\{x_{i}\}_{i=1}^{N}$ with $x_{i}\in\mathbb{R}^{3}$ and $\|x_{i}\|=1$, the pairwise Coulomb energy is defined as
$E=\sum_{i\neq j}\frac{1}{\|x_{i}-x_{j}\|^{2}}.$ (2)
The gradient of this energy function yields repulsive forces between all point pairs which is used to iteratively update the position of the N3N-3 non-fixed points. After each update step, points are re-projected onto the sphere to enforce the unit-norm constraint. By visually enclosing the volume in a sphere, both centered at the origin, we guarantee that all possible slice orientations, which are defined by points on the sphere, will intersect the volume. For each volume, canonical and non-canonical slices are then extracted by first aligning the slicing plane such that its normal vector points to the respective point sampled on the surface of the sphere, and second by identifying the largest lesion slice from the respective canonical and non-canonical plane. Repeating this process for all $i=1,\dots,N$ yields a set of $N$ uniformly distributed omnidirectional slices, thereby effectively resembling the spatial structure of the input volume.
@@
     Compute Coulomb force from all other points $j\neq i$:;
        $F_{i}\leftarrow\sum_{j\neq i}\frac{1}{\|x_{i}-x_{j}\|^{2}}$;
      Update position: $x_{i}\leftarrow x_{i}+\eta\cdot F_{i}$;
@@
Output: Optimized set of $N$ points on the sphere
@@ 3.2 Cross-sectional Image Encoding
After omnidirectional volume slicing, we transfer the obtained 22D volume slices into 11D feature representations by means of DINOv2 [oquab2023dinov2]. Generally, the DINO [caron2021emerging] framework is a self-supervised learning method based on a student–teacher architecture designed to extract semantically rich and meaningful visual representations. Building on this, oquab2023dinov2 [oquab2023dinov2] introduced DINOv2, an extension of DINO [caron2021emerging] trained with self-supervised contrastive learning on a large-scale dataset of 142 million curated images. This large-scale training enables DINOv2 to learn highly transferable semantic features, which have demonstrated strong performance across diverse vision tasks, including both pixel-level tasks (e.g., segmentation and detection) and image-level tasks such as classification. In our framework, we employ a pretrained DINOv2 vision transformer (ViT) model in its small variant, featuring 2121M parameters. The encoder is kept frozen during all experiments and processes input slices of size $224\times 224$ pixels, obtained by linearly up-sampling the input images if not matched initially. The model produces a $384$-dimensional feature representation for a single input image corresponding to a single view.
@@ 3.3 Spherical Mesh-based Graph Topology Construction
Given the $N$ view points, which are equidistantly distributed on the spherical surface (see Section 3.1) and their corresponding cross-sectional volume slice encodings (see Section 3.2), we proceed by describing the construction of the spherical mesh-based graph topology. Therein, nodes represent canonical and non-canonical view planes and slice encodings their corresponding feature vector. We denote the set of points uniformly distributed on the spherical surface as $\mathbf{P}=\{\mathbf{p}_{1},\dots,\mathbf{p}_{N}\}$. To derive a graph structure from the sampled points, we apply Delaunay triangulation to $\mathbf{P}$, which guarantees that no point lies inside the circumcircle of any spherical triangle [delaunay1934sphere]. The triangulation defines the edge set
@@
yielding an undirected graph $\mathbf{G}=(\mathbf{V},\mathbf{E},\mathbf{A})$, where the vertices $\mathbf{V}=\mathbf{P}$ and $\mathbf{A}\in\mathbb{R}^{N\times N}$ is the weighted adjacency matrix, where all connections $\mathbf{e}_{ij}$ resulting from the triangulation have an edge weighting $\mathbf{A}_{ij}=\mathbf{w}_{ij}=1$. This procedure results in a mesh-based graph topology with near-uniform node distribution and well-defined local neighborhoods. Two properties that have been proven to be advantageous for subsequent mesh-based graph learning tasks [bronstein2017geometric]. A graphical visualization of the Delaunay triangulation (step III) and 33D surface mesh extraction (step IV) can be found in Figure 3. Besides the local connectivity pattern, where points $\mathbf{p}_{i}$ and $\mathbf{p}_{j}$ are connected when they are adjacent in the triangulation, we also introduce weighted cross-connections between all node pairs, effectively forming a complete graph. The corresponding edge weights $\mathbf{w}_{ij}$ of cross-connections are defined as the inverse node distance. Here, the distance refers to the hop distance (i.e., shortest path length between two nodes $(\mathbf{v}_{i},\mathbf{v}_{j}\in\mathbf{V}$) using Dijkstra algorithm [dijkstra2022note], defined as
@@ 3.4 Graph Representation Learning
where $\mathbf{W}$ is a learnable weight matrix, $\|$ represents the concatenation operation, and $\mathbf{e}_{ji}$ is the scalar weight on the edge from node $j$ to node $i$. As using multiple node neighborhood aggregation functions has been shown to provide notable improvements [corso2020principal], we leverage both mean and max aggregation. They are particularly suitable to capture the distribution of elements within the graph (i.e., mean) and prove to be advantageous to identify representative elements (i.e, max) [xu2018powerful]. As our underlying graph topology does not change within a setting of $N$ nodes, we do not use sum aggregation, which has been shown to be especially powerful for learning structural graph properties. After the node updates, the resulting set of $N$ one-dimensional node feature vectors is aggregated using dimension-wise mean pooling $\delta$ (i.e., a mean readout function) to obtain a global graph representation $\mathbf{h}_{G}\in\mathbb{R}^{384}$. This representation is then passed through a linear layer $\sigma$ to predict the binary class label $\mathbf{y}_{G}$ for the entire graph, formally expressed as
@@ 4 Experiments & Results
We begin our experimental evaluation section by first describing the datasets used in this study (Section 4.1), followed by implementation and training details (Section 4.2). We then benchmark our proposed volume slicing strategy in Section 4.3 before performing an evaluation on our TomoGraphView framework. We further extend our experimental analysis by investigating the impact of volume slice spacing (Section 4.5), a graph topology ablation study (Section 4.6), and other slice-wise feature aggregation methods (Section 4.7). Finally, we benchmark our proposed TomoGraphView framework against 33D pretrained models (Section 4.8), comparing 33D frozen backbone and 33D model fine-tuning performances.
Figure 4: Visual summary of the collected datasets from various target structures among the human body, including a visual example from each dataset and the number of images under different acquisition criteria. Here, CT refers to computed tomography, MRI refers to magnetic resonance imaging, T1c refers to T1-weighted post-contrast MRI, T2Fs refers to T2-weighted fat-saturated MRI, and DCE refers to dynamic contrast-enhanced MRI.
@@ 4.1 Datasets
In our experimental evaluation, we concentrate on six clinically relevant oncological tasks. A comprehensive overview of the datasets used is provided below, while Figure 4 presents a visual summary of these datasets.
@@ 4.2 Implementation and Training Details
All volumes were first aligned to a common positive Right–Anterior–Superior (RAS+) orientation and subsequently resampled to an isotropic voxel spacing of $1\times 1\times 1$ mm. In order to minimize distortion when resizing slices to match the DINOv2 input resolution, we applied symmetric cropping around the lesion. For our proposed omnidirectional slicing strategy, scalar image volumes were resampled using linear interpolation, while segmentation masks were resampled with nearest-neighbor interpolation to preserve their categorical labels. For all experiments, we attach a lightweight classification head to the frozen DINOv2 feature extractor. Specifically, the head of each comparison method consists of two layers with a non-linear ReLU activation in between, featuring a total of 100100k trainable parameters to ensure a fair comparison across methods. We restrict the prediction head to a shallow design to avoid overfitting, given the limited dataset sizes in medical imaging, which also helps to mitigate graph oversmoothing in the case of GNNs [keriven2022not]. For the MLP, slice encodings are concatenated. We train and evaluate our models using a stratified 5-fold cross-validation scheme according to the binary class label. More specifically, we use three splits for training, one for validation, and one for testing. Training is conducted using the SGD optimizer with a weight decay of $1e^{-3}$, a batch size of $16$, and $300$ epochs. The model that achieves the best AUROC score on the validation set is selected for final evaluation on the test set. A learning rate scheduler first linearly increases the learning rate to $0.001$ for $100$ warm-up epochs, and decays the learning rate afterwards by a factor of $0.95$ when there is no improvement in AUROC validation score for at least 5 epochs, to ensure smooth convergence. All models were trained on an Nvidia RTX A6000. We publicly share our code base [^1] http://github.com/compai-lab/2025-MedIA-kiechle and provide a user-friendly library for omnidirectional volume slicing [^2] https://pypi.org/project/OmniSlicer.
Table 1: Quantitative benchmarking of traditional volume slicing strategies (22D-axial, 22D-axial+, 2.52.5D, 2.52.5D+) against our proposed omnidirectional slicing strategy. All methods are evaluated for various numbers of volume slices, along with an MLP prediction head, using 5-fold cross-validation and AUROC as the performance metric. The rightmost column reports the average performance across datasets. The best-performing method on a per-dataset basis is indicated in bold.
Pred. Head Volume Slicing Strategy Views Breast Brain Head-Neck Kidney Liver Soft-Tissue Mean
MLP 2D-axial 1 0.5844 0.9013 0.6192 0.6541 0.8821 0.7421 0.7305
2D-axial+ 3 0.6166 0.8936 0.6550 0.6234 0.8803 0.7322 0.7335
@@
4.3 Volume Slicing Strategy Benchmarking
In this experiment, we compare the baseline slicing strategies, i.e., 2D-axial, 2D-axial+, 2.5D, and 2.5D+, against our proposed omnidirectional volume slicing method introduced in Section 3.1. To ensure a fair comparison across all datasets and view configurations, each model employs an identical MLP prediction head comprising 100100k trainable parameters. The number of views is empirically set to $8$, $16$, and $24$, providing a reasonable tradeoff between computational efficiency and slice diversity. The corresponding results are summarized in Table 1.

### v2

@@ 3 Methodology
In this work, we present TomoGraphView, a novel framework that aims to preserve the spatial structure of volumetric medical imaging data while leveraging the representational power of a 22D foundation model encoder. Generally, our framework is built upon two key components: omnidirectional volume slicing and spherical mesh–based graph topology construction with corresponding GNN-based feature aggregation and downstream prediction. A visual summary of the overall framework workflow is depicted in Figure 1.
The following sections detail each component of our TomoGraphView framework. We first describe omnidirectional volume slicing in Section 3.1 and relate it to conventional slicing approaches. Section 3.2 outlines the encoding of the obtained omnidirectional 22D cross-sectional slices into 11D feature representations. In Section 3.3, we introduce the spherical mesh–based graph topology construction, followed by Section 3.4, which describes the graph representation learning process used for downstream classification.
@@ 3.1 Omnidirectional Volume Slicing
Figure 2: Visual representation of different volume slicing strategies and resulting slices for a synthetically generated input volume depicting an elongated shape diagonally within the volume (i.e., not aligned with any canonical axis). The presented volume slicing strategies include 22D-axial and 22D-axial+ (left), 2.52.5D and 2.52.5D+ (middle), and our proposed omnidirectional volume slicing (right). Green slices represent slices extracted from the axial direction, red from the sagittal direction, and blue from the coronal direction, whereas yellow shows out-of-plane slices as obtained by means of our omnidirectional volume slicing strategy.
@@
In this work, additionally to utilizing information from canonical planes, we propose to extract slices from non-canonical orientations to obtain omnidirectional slice representations, thereby enabling a richer characterization of volumetric structures that are not naturally aligned with the canonical axes. To obtain $N$ omnidirectional slice representations, we visually embed the volume within a bounding sphere $\mathcal{S}$, centered at the origin, which is formally described as
@@
where $r$ is the radius such that the sphere fully contains the 33D volumetric scan as illustrated in Figure 3. Next, as we intend to make use of both, canonical and non-canonical perspectives, we first of all fix three points on the surface of the sphere representing the canonical view planes (i.e., axial, coronal, and sagittal) with their known coordinates $(r,0,0),(0,r,0)$ and $(0,0,r)$, as illustrated in Figure 3 (step I). We then optimize the location of the remaining (N3)(N-3) points, representing non-canonical view planes, such that all points on the surface are approximately equally distributed, as illustrated in Figure 3 (step II). To this end, we make use of the repulsion-based optimization outlined in Algorithm 1, which is inspired by the Thomson problem [thomson1904xxiv]. Therein, points are interpreted as identical charged particles constrained to lie on the spherical surface, and their positions are optimized by minimizing the system’s total electrostatic energy. Formally, given $N$ points $\{x_{i}\}_{i=1}^{N}$ with $x_{i}\in\mathbb{R}^{3}$ and $\|x_{i}\|=1$, the pairwise simplified Coulomb energy, assuming unit charges and ignoring the Coulomb constant, is used without loss of generality, and defined as
$E(N)=\sum_{i\neq j}\frac{1}{\|x_{i}-x_{j}\|^{2}}.$ (2)
The gradient of the Coulomb energy function $\nabla E(N)$ yields repulsive forces between all point pairs which is used to iteratively update the position of the N3N-3 non-fixed points using gradient descent. After each update step, points are re-projected onto the sphere to enforce the unit-norm constraint. By visually enclosing the volume in a sphere, both centered at the origin, we guarantee that all possible slice orientations, which are defined by points on the sphere, will intersect the volume. For each volume, canonical and non-canonical slices are then extracted by first aligning the slicing plane such that its normal vector points to the respective point sampled on the surface of the sphere, and second by identifying the largest lesion slice from the respective canonical and non-canonical plane. Repeating this process for all $i=1,\dots,N$ yields a set of $N$ uniformly distributed omnidirectional slices, thereby effectively resembling the spatial structure of the input volume.
@@
     Compute Coulomb force between pairs $j\neq i$;
        $E_{i}\leftarrow\sum_{j\neq i}\frac{1}{\|x_{i}-x_{j}\|^{2}}$;
      Update position: $x_{i}\leftarrow x_{i}-\eta\cdot\nabla E_{i}$;
@@
0.8em Output: Optimized set of $N$ points on the sphere
0.8em
@@ 3.2 Cross-sectional Image Encoding
After omnidirectional volume slicing, we transfer the obtained 22D volume slices into 11D feature representations by means of DINOv2 [oquab2023dinov2]. Generally, the DINO [caron2021emerging] framework is a self-supervised learning method based on a student–teacher architecture designed to extract semantically rich and meaningful visual representations. Building on this, oquab2023dinov2 [oquab2023dinov2] introduced DINOv2, an extension of DINO [caron2021emerging] trained with self-supervised contrastive learning on a large-scale dataset of 142 million curated images. In this work, we employ a pretrained DINOv2 vision transformer (ViT) model in its small variant, featuring 2121M parameters. The encoder is kept frozen during all experiments and processes input slices of size $224\times 224$ pixels, obtained by linearly upscaling the input images if they are not initially matched. The model produces a $384$-dimensional feature representation for a single input image corresponding to a single view.
@@ 3.3 Spherical Mesh-based Graph Topology Construction
Given the $N$ view points, which are equidistantly distributed on the spherical surface (Section 3.1) and their corresponding cross-sectional volume slice encodings (Section 3.2), we proceed by describing the construction of the spherical mesh-based graph topology. Therein, nodes represent canonical and non-canonical view planes and slice encodings their corresponding feature vector. We denote the set of points uniformly distributed on the spherical surface as $\mathbf{P}=\{\mathbf{p}_{1},\dots,\mathbf{p}_{N}\}$. To derive a graph structure from the sampled points, we apply Delaunay triangulation to $\mathbf{P}$, which guarantees that no point lies inside the circumcircle of any spherical triangle [delaunay1934sphere]. The triangulation defines the edge set
@@
yielding an undirected graph $\mathbf{G}=(\mathbf{V},\mathbf{E},\mathbf{A})$, where the vertices $\mathbf{V}=\mathbf{P}$ and $\mathbf{A}\in\mathbb{R}^{N\times N}$ is the weighted adjacency matrix, where all connections $\mathbf{e}_{ij}$ resulting from the triangulation have an edge weighting $\mathbf{A}_{ij}=\mathbf{w}_{ij}=1$. This procedure results in a mesh-based graph topology with near-uniform node distribution and well-defined local neighborhoods. Two properties that have been proven to be advantageous for subsequent mesh-based graph learning tasks [bronstein2017geometric]. A graphical visualization of the Delaunay triangulation (step III) and 33D surface mesh extraction (step IV) can be found in Figure 3. Besides the local connectivity pattern, where points $\mathbf{p}_{i}$ and $\mathbf{p}_{j}$ are connected when they are adjacent in the triangulation, we also introduce weighted cross-connections between all remaining node pairs, effectively forming a complete graph. The corresponding edge weights $\mathbf{w}_{ij}$ of cross-connections are defined as the inverse node distance. Here, the distance refers to the hop distance (i.e., shortest path length between two nodes $(\mathbf{v}_{i},\mathbf{v}_{j}\in\mathbf{V}$) using Dijkstra algorithm [dijkstra2022note], defined as
@@ 3.4 Graph Representation Learning
where $\mathbf{W}$ is a learnable weight matrix, $\|$ represents the concatenation operation, and $\mathbf{e}_{ji}$ is the scalar weight on the edge from node $j$ to node $i$. As using multiple node neighborhood aggregation functions has been shown to provide notable improvements [corso2020principal], we leverage both mean and max aggregation. They are particularly suitable to capture the distribution of elements within the graph (i.e., mean) and prove to be advantageous to identify representative elements (i.e, max) [xu2018powerful]. As our underlying graph topology does not change within a setting of $N$ nodes, we do not use sum aggregation, which has been shown to be especially powerful for learning structural graph properties. After the node updates, the resulting set of $N$ one-dimensional node feature vectors is aggregated using dimension-wise mean pooling $\delta$ (i.e., a mean readout function) to obtain a global graph representation $\mathbf{h}_{G}\in\mathbb{R}^{N}$. This representation is then passed through a linear layer $\sigma$ to predict the binary class label $\mathbf{y}_{G}$ for the entire graph, formally expressed as
@@
Figure 4: Visual summary of the collected datasets from various target anatomical structures across the human body. The overview includes a visual example from each dataset (left), as well as the number of images under different acquisition criteria (right). Here, CT refers to computed tomography images, MRI refers to magnetic resonance imaging, T1c refers to T1-weighted post-contrast, T2Fs refers to T2-weighted fat-saturated, and DCE refers to dynamic contrast-enhanced.
@@ 4 Experiments & Results
Our experimental evaluation begins with a description of the datasets used in this study (Section 4.1), followed by details on implementation and training procedures (Section 4.2). We then assess our proposed volume-slicing strategy (Section 4.3), before evaluating the full TomoGraphView framework (Section 4.4). The analysis is extended by examining the effect of slice spacing (Section 4.5), conducting a graph-topology ablation (Section 4.6), and comparing alternative slice-wise feature aggregation methods (Section 4.7). Finally, we benchmark TomoGraphView against 33D pretrained models (Section 4.8), considering both frozen and finetuned 33D backbone configurations.
@@ 4.1 Datasets
We focus our experimental evaluation on six clinically meaningful oncological tasks. An overview of the datasets used is summarized below, and Figure 4 provides a visual depiction.
@@ 4.2 Implementation and Training Details
All volumes were first aligned to a common positive Right–Anterior–Superior (RAS+) orientation and subsequently resampled to an isotropic voxel spacing of $1\times 1\times 1$ mm. In order to minimize distortion when resizing slices to match the DINOv2 input resolution, we applied symmetric cropping around the lesion. For our proposed omnidirectional slicing strategy, scalar image volumes were resampled using linear interpolation, while segmentation masks were resampled with nearest-neighbor interpolation to preserve their categorical labels. For all experiments, we attach a lightweight classification head to the frozen DINOv2 feature extractor. Specifically, the head of each baseline method consists of two layers with a non-linear ReLU activation in between, featuring a total of 100100k trainable parameters to ensure a fair comparison across methods. We restrict the prediction head to a shallow design to avoid overfitting, given the limited dataset sizes in medical imaging, which also helps to mitigate graph oversmoothing in the case of GNNs [keriven2022not]. For the MLP, slice encodings are concatenated. We train and evaluate our models using a 5-fold cross-validation scheme, and stratify splits according to the binary class label. More specifically, we use three splits for training, one for validation, and one for testing. Training is conducted using the SGD optimizer with a weight decay of $1e^{-3}$, a batch size of $16$, and $300$ epochs. The model that achieves the best AUROC score on the validation set is selected for final evaluation on the test set. A learning rate scheduler first linearly increases the learning rate to $0.001$ for $100$ warm-up epochs, and decays the learning rate afterwards by a factor of $0.95$ when there is no improvement in AUROC validation score for at least 5 epochs, to ensure smooth convergence. All models were trained on an Nvidia RTX A6000 GPU. We publicly share our code repository [^1] http://github.com/compai-lab/2025-MedIA-kiechle and furthermore provide a user-friendly library for omnidirectional volume slicing [^2] https://pypi.org/project/OmniSlicer.
Table 1: Quantitative comparison of conventional volume slicing strategies (2D-axial, 2D-axial+, 2.5D, 2.5D+) with our proposed omnidirectional slicing approach using an MLP prediction head, as well as our TomoGraphView framework. All methods are evaluated across multiple datasets and numbers of volume slices (views) using 5-fold cross-validation and Area Under the Receiver Operating Characteristic Curve (AUROC) as the performance metric. The rightmost column reports the mean performance over all datasets. Bold values indicate the best result per dataset, and underlined values denote the second-best.
Method (Head – Slicing Strategy) Views Breast Brain Head-Neck Kidney Liver Soft-Tissue Mean
MLP – 2D-axial Slicing 1 0.5844 0.9013 0.6192 0.6541 0.8821 0.7421 0.7305
MLP – 2D-axial+ Slicing 3 0.6166 0.8936 0.6550 0.6234 0.8803 0.7322 0.7335
@@
8 0.7336 0.9678 0.6780 0.7878 0.9600 0.8399 0.8279
16 0.6907 0.9666 0.7244 0.7479 0.9585 0.8389 0.8212
GNN – Omnidirectional Slicing TomoGraphView (ours) 24 0.7084 0.9655 0.7527 0.7893 0.9634 0.8441 0.8372
4.3 Impact of Volume Slicing Strategy
We start our experimental evaluation by comparing baseline volumetric slicing strategies, i.e., 2D-axial, 2D-axial+, 2.5D, and 2.5D+, against our proposed omnidirectional volume slicing method introduced in Section 3.1. To ensure a fair comparison across all datasets and view configurations, our volume slicing strategy benchmark is solely based on identical MLP prediction heads comprising 100100k trainable parameters. The number of views is empirically set to $8$, $16$, and $24$, providing a reasonable tradeoff between computational efficiency and captured slice diversity. The corresponding results for any combination of MLP and the respective volume slicing strategy are summarized in Table 1, and refer to all non-gray-colored rows.